Ticket Number: HEA-NOC/20020425-01 Ticket Status: CLOSE
Ticket Type: Unscheduled Ticket Source: HEA-NOC
Ticket Scope: Site Site/Line: HEAnet <==> NCIR
Ticket Owner: [log in to unmask] Problem Fixer: HEA-NOC
Ticket Opened: Problem Start: 20020425 09:33
Ticket Closed: 20020515 11:12:42 UTC Problem End:
Brief loss of connection between HEAnet and NCIR
20020426 12:12 Circuit goes down again.
All traffic between HEAnet and NCIR
09:32 Loss of connectivity between HEAnet and NCI.
09:33 Route between sites returns.
09:45 AB contacts NCI. No reported problems on site.
09:55 AB contacts Eicom and requests alarm check on line. Ticket 2212 issued by Eircom.
10:41 Eircom report no errors on line. Ticket 2212 closed by Eircom.
20020426 12:12 Loss of connectivity between HEAnet and NCI.
20020426 12:18 AB contacts NCI to report loss of circuit. AB leaves voicemail outlining problem.
20020426 12:22 AB contacts Eircom requesting alarm check. Ticket 2359 opened.
20020426 12:50 NCIR report that the outage is as a result of a power surge locally. They are currently working to get all services back online. Offsite MX backup is at Kerna.
20020426 13:40, 14:18 NCI call to confirm that the power is still down.
20020426 15:00 NCI call and state that modem on-site broke as a result of the power surge. Eircom are currenlty replacing the modem with a new device. It is not yet known the time to fix.
20020426 15:17 NTU replaced. Line/ Protocol comes back up. Traffic flows as normal.
20020503 12:00 UTC+1 CW line reset at 0918 - unable to contact NCIRL
20020503 1233 CW NCIRL report no issues on site reporting to eircom ticket 288 Requesting an alarm check for 0900-1000hrs
20020503 1440 CW Eircom no errors detected.
20020515 112 CW Closing ticket, no further issues on line
Time to Fix:
1 min + 3hours Additional Notes:
20020426 15:18 Ticket remains open until circuit is stable.